The Botosan Factory has determined that its budgeted factory overhead budget for the year is $696,864 and budgeted direct labor hours are 292,800. If the actual direct labor hours for the period are 266,400, how much overhead would be allocated to the period?

a.$634,032

b.$538,927

c.$653,053

d.$767,179
Answer #1

Overhead rate =Overhead budget/direct labor hours =696864/292800 =2.38

Overhead allocated to the period =266400*2.38 =634032